OpenStudy (anonymous):

Someone please help me!!!! Im so desperate What is the simplified form of the quantity 1 over y minus 1 over x all over 1 over y plus 1 over x ?

OpenStudy (anonymous):

@campbell_st can you help me at all? If not can you help with another you may help with?

OpenStudy (campbell_st):

ok... so is the question \[\frac{\frac{\frac{1}{y -1}}{x}}{\frac{\frac{1 }{y + 1}}{x}}\]

OpenStudy (anonymous):